# On the Identifying Power of Generalized Monotonicity

### v3

@@ Appendix B Details of Examples
The condition in (35) states that if crossing the cutoff for field 1 does not cause the student to switch to field 1, then it does not cause them to switch to or away from field 2. A similar interpretation applies to (36). Lee and Salanié (2023) show the set of all distributions that satisfy (33)–(36) are equivalent to a strict one-to-one targeting model with $|\mathcal{Z}|=3$ and $|\mathcal{D}^{\dagger}|=2$; it therefore follows from Remark 4.5 that any $Q$ that satisfies (33)–(36) also satisfies Assumption 2.2. Here, we establish directly that (33)–(36) imply Assumption 2.2 with $z^{\ast}(0)=0$, $z^{\ast}(1)=1$, and $z^{\ast}(2)=2$. To show $z^{\ast}(0)=0$, we prove by contradiction that

### v4

@@ Appendix B Details of Examples
The condition in (36) states that if crossing the cutoff for field 1 does not cause the student to switch to field 1, then it does not cause them to switch to or away from field 2. A similar interpretation applies to (37). lee2023treatment show the set of all distributions that satisfy (34)–(37) are equivalent to a strict one-to-one targeting model with $|\mathcal{Z}|=3$ and $|\mathcal{D}^{\dagger}|=2$; it therefore follows from Remark 4.5 that any $Q$ that satisfies (34)–(37) also satisfies Assumption 2.2. Here, we establish directly that (34)–(37) imply Assumption 2.2 with $z^{\ast}(0)=0$, $z^{\ast}(1)=1$, and $z^{\ast}(2)=2$. To show $z^{\ast}(0)=0$, we prove by contradiction that